# Solving the Tree Containment Problem Using Graph Neural Networks

### v1

@@ 1 Introduction
    This work shows the potential of using GNNs for solving phylogenetic problems. Potentially, similar approaches might be applied to solving different problems.
@@ 4.1 The summary of our approach
After obtaining a network-tree graph, we apply a GNN to it. The architecture of the used GNN is depicted in Figure 2 (right side). The goal of the GNN is to extract the necessary information from the graph structure and produce the answer to the tree containment problem. First, we use a multi-layer message-passing GNN to obtain meaningful embeddings of the nodes (comprehensive survey of various GNNs can be found, for instance, in Hamilton (2020)).
@@ 5.1 Data
We vary the number of leaves in the generated datasets. We consider two different scenarios: 1) Transductive learning: the number of leaves in the validation/test instances does not exceed the number of leaves in the training data; 2) Inductive learning: the number of leaves in the validation/test instances is in the interval $(L_{train};2L_{train}]$, while the number of leaves in the training set is in $[L_{min},L_{train}]$. The minimum number of leaves $L_{min}$ in all instances is 5, and the maximum number considered is 100. The summary statistics of the generated graphs are provided in Appendix, Figure 10.
@@
For the the real-world experiments, we use the dataset of gene trees from bacterial and archaeal genomes (proposed in Beiko (2011), binarized in van Iersel et al. (2022)). For each set of input trees, we construct a network with the minimized number of reticulations such that it contains all given trees using the TrivialRand algorithm from Bernardini et al. (2023). Then, to obtain negative examples, we perform tail moves on the network similar to the synthetic datasets construction. As obtaining the correct tree containment labels using BOTCH is becoming extremely time consuming for large networks (that have large number of leaves and/or high number of reticulations), we limit the considered instances to maximum 50 leaves and 5 input trees. As we have limited number of real-world instances (less than 300 network-tree pairs for each value of $L\in\{10,20,30,40,50\}$, we use them only as validation and test data while using the synthetic LGT data for training. We also do not specifically tune any hyperparameters on the real-world data. For this experiment, we sample the parameter $\alpha$ of LGT generator uniformly randomly from $[0.3;0.5]$ (to increase the number of reticulation nodes in the networks). The number of leaves in the train instances is from $[5;50]$ for all test instances.
@@ 5.2 Baselines
The first baseline we consider is a naive approach that disregards the labels of the leaves. Instead, it attempts to predict the tree containment property based on manually generated features from both the network and the tree, such as the average distance from the root to the leaves. The complete list of used features is listed in Appendix, Table 2. We train an XGBoost model (Chen & Guestrin, 2016) on the dataset using these features.
Next, we consider the Siamese GNN (similar to the GREED algorithm (Ranjan et al., 2022)) applied to the tree containment problem. We make it aware of the labels of the leaves by using one-hot encoding of the labels as the node features. This approach does not enable the inductive learning ability, but could potentially work well in the transductive learning setting. Moreover, in theory, it can use the automatically extracted features from the network and the tree (in contrast to the manually constructed features used in the first considered baseline).
@@ 5.3 GNN architecture and training hyperparameters tuning
We performed a grid search to find the best performing architecture and the corresponding training hyperparameters of the GNN in the Combine-GNN. The tuned hyperparameters and their ranges are the following (the best values are underlined): initial learning rate: {$10^{-4},\underline{10^{-3}},10^{-2}$}; weight decay: $\underline{0},10^{-4},10^{-3}$}; dropout: {$0.0,\underline{0.2}$}; number of GNN layers: {$3,\underline{5}$}, size of node embeddings: {$32,64,\underline{128}$}; type of GNN operation: {GCN, GIN, GAT}; readout operation: {mean, max, sum}. Grid search hyperparameter tuning was also performed for the baselines. The best hyperparameters for them are provided in Appendix, Table 1. During hyperparameter tuning, we used validation subsets and random seeds different from those used to report our main results.
@@ 6 Results
The performance of Combine-GNN in test datasets with instances of increasing size (including those which are much larger than the instances contained in the corresponding training dataset) is presented in Figure 4. We observe that while the performance deteriorates with the difference between size of the training and test instances increasing, Combine-GNN demonstrates rather robust behaviour, for instance, when trained on instances with up to 20 leaves, the balanced accuracy score on datasets with instances with up to 100 leaves, stays above 0.9. This highlights the generalization ability of the proposed approach.
@@
Figure 6: Performance (balanced accuracy) for training and testing datasets with varying percentage of reticulation nodes. Here, the transductive setup denotes the test instances with $L$ from $[5;10]$ and inductive denotes the instances with $L$ from $[11;20]$. The results are averaged over five runs with different seeds for each setup and each percentage of reticulation nodes.
The performance of Combine-GNN with training and testing datasets of varying number of reticulation nodes is presented in Figure 6. We use as training datasets the datasets with instances with $[5;10]$ leaves and varying number of reticulations (by controlling the parameter $\alpha$ in the LGT generator). The test datasets have been generated using the same $\alpha$ correspondingly. For a smaller number of samples, the number of epochs has been adjusted accordingly, so that the total number of processed samples during training is equal in all cases ($10^{6}$). We can conclude that a larger percentage (number) of reticulations makes the task much more difficult for Combine-GNN to solve. However, we note that in the transductive setup, the accuracy stays above 0.9 even for the dataset with >20% reticulation nodes.
@@
The results for real-world data experiments are shown in Figure 7. We note the challenges posed in this experiment: the training data is a synthetic dataset generated using LGT generator (real-world data is used only as validation/test subsets); furthermore, the used hyperparameters were tuned on synthetic datasets only; the real-world validation and test subsets are rather small causing potential overfitting while we choose the best epoch based on the validation performance.

### v2

@@ 1 Introduction
    This work shows the potential of using GNNs for solving phylogenetic problems. Potentially, ideas from the proposed approach to solving tree containment, might be applied to solving other phylogenetic problems as well.
@@ 4.1 The summary of our approach
After obtaining a network-tree graph, we apply a GNN to it. The architecture of the used GNN is depicted in Figure 2 (right side). The goal of the GNN is to extract the necessary information from the graph structure and produce the answer to the tree containment problem. First, we use a multi-layer message-passing GNN to obtain meaningful embeddings of the nodes (a comprehensive survey of various GNNs can be found, for instance, in Hamilton (2020)).
@@ 5.1 Data
We vary the number of leaves in the generated datasets. We consider two different scenarios: 1) Transductive learning: the number of leaves in the validation/test instances does not exceed the number of leaves in the training data; 2) Inductive learning: the number of leaves in the validation/test instances is in the interval $(L_{train};2L_{train}]$, while the number of leaves in the training set is in $[L_{min},L_{train}]$. The minimum number of leaves $L_{min}$ in all instances is 5, and the maximum number considered is 100. We sample the parameter $\alpha$ of the LGT generator uniformly randomly from $[0.1;0.2]$. The summary statistics of the generated graphs are provided in Appendix, Figure 10.
@@
For the the real-world experiments, we use the dataset of gene trees from bacterial and archaeal genomes (proposed in Beiko (2011), binarized in van Iersel et al. (2022)). For each set of input trees, we construct a network with the minimized number of reticulations such that it contains all given trees using the TrivialRand algorithm from Bernardini et al. (2023). Then, to obtain negative examples, we perform tail moves on the network similar to the synthetic datasets construction. As obtaining the correct tree containment labels using BOTCH is becoming extremely time consuming for large networks (that have large number of leaves and/or high number of reticulations), we limit the considered instances to maximum 50 leaves and 5 input trees. As we have limited number of real-world instances (less than 300 network-tree pairs for each value of $L\in\{10,20,30,40,50\}$, we use them only as validation and test data while using the synthetic LGT data for training. We also do not specifically tune any hyperparameters on the real-world data. For this experiment, we sample the parameter $\alpha$ of the LGT generator uniformly randomly from $[0.3;0.5]$ (to increase the number of reticulation nodes in the networks). The number of leaves in the train instances is from $[5;50]$ for all test instances.
@@ 5.2 Baselines
The first baseline we consider is a naive approach that disregards the labels of the leaves. Instead, it attempts to predict the tree containment property based on manually generated features from both the network and the tree, such as the average distance from the root to the leaves. The complete list of used features is listed in Appendix, Table 2. We train an XGBoost model (Chen & Guestrin, 2016) using these features.
Next, we consider a Siamese GNN (similar to the GREED algorithm (Ranjan et al., 2022)) applied to the tree containment problem. We make it aware of the labels of the leaves by using one-hot encoding of the labels as the node features. This approach does not enable the inductive learning ability, but could potentially work well in the transductive learning setting. Moreover, in theory, it can use the automatically extracted features from the network and the tree (in contrast to the manually constructed features used in the first considered baseline).
@@ 5.3 GNN architecture and training hyperparameters tuning
We performed a grid search to find the best performing architecture and the corresponding training hyperparameters for Combine-GNN. The tuned hyperparameters and their ranges are the following (the best values are underlined): initial learning rate: {$10^{-4},\underline{10^{-3}},10^{-2}$}; weight decay: $\underline{0},10^{-4},10^{-3}$}; dropout: {$0.0,\underline{0.2}$}; number of GNN layers: {$3,\underline{5}$}, size of node embeddings: {$32,64,\underline{128}$}; type of GNN operation: {GCN, GIN, GAT}; readout operation: {mean, max, sum}. Grid search hyperparameter tuning was also performed for the baselines. The best hyperparameters for them are provided in Appendix, Table 1. During hyperparameter tuning, we used validation subsets and random seeds different from those used to report our main results.
@@ 6 Results
The performance of Combine-GNN on test datasets with instances of increasing size (including those which are much larger than the instances contained in the corresponding training dataset) is presented in Figure 4. We observe that while the performance deteriorates with the difference between size of the training and test instances increasing, Combine-GNN demonstrates rather robust behaviour, for instance, when trained on instances with up to 20 leaves, the balanced accuracy score on datasets with instances with up to 100 leaves, stays above 0.9. This highlights the generalization ability of the proposed approach.
@@
Figure 6: Performance (balanced accuracy) for training and testing datasets with varying percentage of reticulation nodes. Here, the transductive setup denotes the test instances with $L$ from $[5;10]$ and inductive denotes the instances with $L$ from $[11;20]$. The training instances in both setups have the $L$ from with $[5;10]$. The results are averaged over five runs with different seeds for each setup and each percentage of reticulation nodes.
The performance of Combine-GNN with training and testing datasets of varying number of reticulation nodes is presented in Figure 6. We use as training datasets the datasets with instances with $[5;10]$ leaves and varying number of reticulations (by controlling the parameter $\alpha$ in the LGT generator). The test datasets have been generated using the same $\alpha$ correspondingly. We can conclude that a larger percentage (number) of reticulations makes the task much more difficult for Combine-GNN to solve. However, we note that in the transductive setup, the accuracy stays above 0.9 even for the dataset with >20% reticulation nodes.
@@
The results for real-world data experiments are shown in Figure 7. We note the challenges posed in this experiment: the training data is a synthetic dataset generated using LGT generator (real-world data is used only as validation/test subsets); furthermore, the used hyperparameters were tuned on synthetic datasets only; the real-world validation subsets are rather small causing potential overfitting on them (we choose the best epoch based on the validation performance).
@@ 8 Conclusion
Acknowledgements
The work in this paper is supported by the Dutch Research Council (NWO) through project OCENW.GROOT.2019.015 "Optimization for and with Machine Learning (OPTIMAL)".